A small ball is released from rest and falls on the horizontal platform which is
descending with a constant speed 7 1ms-1 .Given that the ball is 12m above the
platform at the instant of release. Calculate the time that elapses before the ball
hits the plat form.
1
Expert's answer
2020-04-09T09:29:33-0400

As per the given question,

Initial height from the floor of the platform "(h_1)=12m"

Speed of the platform "(v)=71 m\/sec"

Let after time t, both will be at the same height,

Hence, distance covered by the platform in t time is x,

"x=71\times t"

Now, height covered by the ball in the same time t will be,

"12+x=ut+\dfrac{gt^2}{2}"

here the initial velocity of the ball is u and g is the gravitational acceleration,

"g= 9.8 m\/sec^2"

now substituting the values in the above,

"12+71t=0+\dfrac{9.8\times t^2}{2}"

"\Rightarrow 12+71t=4.9t^2"

"\Rightarrow 4.9t^2-71t-12=0"

we know that, if "ax^2+bx+c=0"

"x=\dfrac{-b\pm\sqrt{b^2-4ac}}{2a}"

Hence

"t=\dfrac{71\pm\sqrt{71^2+4\times 4.9\times12}}{2\times 4.9}"


"t=\dfrac{71\pm 72.63}{9.8}"

We know that time, never can be negative, hence leaving the negative value.

"t=\dfrac{71+72.63}{9.8}=14.65 sec"